Write the equation of the line through (1,2) and (-3,6) in the y=mx+b format
Answer by stanbon(75887)

Write the equation of the line through (1,2) and (-3,6) in the y=mx+b format
-----------
slope = [6-2]/[-3-1] = 4/-4 = -1
--------------
intercept: 2 = -1(1) + b
b = 3
==========
EQUATION:
y = -x + 3
===============
Cheers,
Stan H.